OpenStudy (anonymous):

Consider the limit: lim_{i rightarrow 1}sum_{i=1}^{n}3(1+4i divn)(4 divn) (a)Evaluate the limit (b) Does the limit define an integral?

OpenStudy (anonymous):

draw it

OpenStudy (anonymous):

draw?

OpenStudy (zarkon):

I assume you mean \[\lim_{n\to\infty}\sum_{i=1}^{n}3\left(1+\frac{4i}{n}\right)\frac{4}{n}\]

OpenStudy (anonymous):

\[3\sum_{i=1}^{n}(16i divn ^{2})\]